C = 450 + 0.4Y
I = 350

G = 150
X = 70
Z = 35 + 0.1Y
T = 0.15Y
Yf = 1550




Q.2.1 Calculate the level of autonomous spending in this economy. (2)
1
Expert's answer
2020-12-03T10:54:59-0500

AE=C+I+G+X+Z+T+Yf

AE=450+0.4Y+350+150+70+35+0.1Y+0.15Y+1550

=2535+0.65Y